1. Chuck, a single taxpayer, earns $70,000 in taxable income and $10,000 in interest from an investment in City of Heflin bonds. Using the U.S. tax rate schedule, how much federal tax will he owe?

2.Chuck, a single taxpayer, earns $70,000 in taxable income and $10,000 in interest from an investment in City of Heflin bonds. What is his average tax rate? (Carry your answer two decimals, i.e., 20.05)

3. Chuck, a single taxpayer, earns $70,000 in taxable income and $10,000 in interest from an investment in City of Heflin bonds. What is his effective tax rate?

4. Chuck, a single taxpayer, earns $70,000 in taxable income and $10,000 in interest from an investment in City of Heflin bonds. What is his current marginal tax rate?
Answer #1

1. Taxable income = $70,000

Chunk will owe federal tax =$987.5+$3629.88+(22% of 29875)$6572.5 = $11,189.88

2. Average rate of tax = Total tax /total income *100 = $11,189.88/$70,000*100 = 15.98%

3. Effective rate of tax = $11,189.88/($70000+10000)=$11,189.88/$80,000*100 = 13.98%

4. Marginal tax rate is slab under which income is taxed which is 22%

Note: US tax rates applicable to 2020-2021 are considered.

Interest from investment in City Heflin bonds are not taxable.